The average age of Samir's parents is 49. His father is 8 years older than his mother. If the average age of Samir and his father is 27, how old is Samir?

Answer by richwmiller(9135)

(m+f)/2=49
f=m+8
(s+f)/2=27
f=53
m=45
s=1